# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

### v1

@@ Appendix A Free-electron self-energy operator
$D_{\mu\nu}^{F}(\mathrm{k})=-\frac{g_{\mu\nu}}{\mathrm{k}^{2}}$ (48)
@@
$D_{\mu\nu}^{C}(\mathrm{k})=\frac{1}{\mathrm{k}^{2}}(-g_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{\mathbf{k}^{2}}+\frac{k^{0}(k_{\mu}\delta_{\nu 0}+k_{\nu}\delta_{\mu 0})}{\mathbf{k}^{2}})$ (49)
in the Coulomb gauge. In Eq. 49, $\delta_{\nu\mu}$ is the four-dimensional Kronecker delta. The expression (47) is UV divergent in both gauges. To eliminate the UV divergences, we apply the renormalization procedure. To make sense of divergent expressions, we use the dimensional regularization ($D=4-2\varepsilon$). In both gauges, we separate out only the UV divergent part by writing the self-energy in the form:
@@
Note that we use this definition instead of writing a more familiar expression in terms of the renormalization constant $Z_{2}$. In the Feynman gauge, the second approach leads to the IR singularities of individual contributions. In the Coulomb gauge, both approaches are equally convenient, but, for consistency, we write the expression for the free-electron self-energy operator in this form. The notation $\Delta_{\varepsilon}=\frac{1}{\varepsilon}-\gamma_{E}+\ln 4\pi+\ln m$ is introduced, and the mass counterterm is
$\delta m=\frac{3\alpha}{4\pi}m(\Delta_{\varepsilon}+\frac{4}{3}).$ (51)
@@
$).\Sigma^{(0)}_{R}(\mathrm{p})=\frac{\alpha}{4\pi}(a(p_{0},p)+\not{\mathrm{p}}b(p_{0},p)+\gamma^{0}c(p_{0}$,p)). (52)
The coefficients $a$, $b$ and $c$ are gauge-dependent. In the Feynman gauge, they can be found, e.g., in Ref. base_1999:
$\begin{gathered}a(p_{0},p)=2m\left(1+\frac{2\rho}{1-\rho}\ln\rho\right),\\ b(p_{0},p)=-\frac{2-\rho}{1-\rho}\left(1+\frac{\rho}{1-\rho}\ln\rho\right),\\ c(p_{0},p)=0,\end{gathered}$ (53)
where $\rho=1-{\mathrm{p^{2}}}/{m^{2}}$. For the Coulomb gauge, they could be derived from the expression base_adkins of the form:
@@
$\begin{gathered}X=1+\frac{p^{2}}{m^{2}}(1-x),\\ Y=1-\frac{\mathrm{p}^{2}}{m^{2}}(1-x)-i0,\\ Z=1-\frac{p_{0}^{2}}{m^{2}}(1-u)+\frac{p^{2}}{m^{2}}(1-xu)-i0.\end{gathered}$ (55)
Some integrals in this expression can be evaluated analytically, which was done, e.g., in Ref. zp_final. The latter results allows one to write for the coefficients $a$, $b$, and $c$:
$\begin{gathered}a(p_{0},p)=2m\left(1-F_{0}+\frac{\rho\log\rho}{1-\rho}\right),\\ b(p_{0},p)=\frac{(\rho-2)(1-\rho+\rho\ln\rho)}{(1-\rho)^{2}}-2F_{2}\rho+\frac{2m^{2}(F_{1}\rho\ln\rho-F_{0})}{p^{2}},\\ c(p_{0},p)=2\frac{p_{0}}{p^{2}}(F_{0}m^{2}-F_{1}m^{2}\rho\ln\rho+F_{2}\rho p^{2}),\end{gathered}$ (56)
where $F_{0}$, $F_{1}$ and $F_{2}$ are
$\begin{gathered}F_{0}=\frac{\sqrt{p^{2}+m^{2}}}{p}\ln\left|\frac{\sqrt{p^{2}+m^{2}}+p}{\sqrt{p^{2}+m^{2}}-p}\right|-2,\\ F_{1}=\frac{p_{0}}{p}\ln\left|\frac{p_{0}+p}{p_{0}-p}\right|-2,\\ F_{2}=\int_{0}^{1}\frac{\sqrt{x}\ln Xdx}{X-\rho+i0}.\end{gathered}$ (57)
@@ Appendix B Free-electron vertex operator
In order to calculate the one-potential term, one needs the free-electron vertex function, given by:
@@
This expression is UV divergent in both gauges. After regularization, it can be written as:
@@
where, as in the case of the free-electron self-energy operator, the UV finite part of $\Gamma_{R}$ also turns out to be IR finite. For brevity, we give below only the time component of the operator, $\Gamma^{0}_{R}$, for the case when $p_{0}=p^{\prime}_{0}=\varepsilon_{a}$:
@@
$\begin{gathered}A=C_{5}-2+\mathrm{p}^{\prime 2}C_{11}+\mathrm{p}^{2}C_{12}+4(\mathrm{p}^{\prime}\cdot\mathrm{p})(C_{00}+C_{11}+C_{12})+m^{2}(-2C_{00}+C_{11}+C_{12}),\\ B_{1}=-4(C_{11}+C_{23}),\;B_{2}=-4(C_{00}+C_{11}+C_{12}+C_{25}),\\ C_{1}=B_{2},\;C_{2}=-4(C_{12}+C_{24}),\\ D=2(C_{00}+C_{11}+C_{12}),\\ H_{1}=4m(C_{00}+2C_{11}),\;H_{2}=4m(C_{00}+2C_{12}),\;G_{1}=G_{2}=0,\end{gathered}$ (61)
@@
$\begin{gathered}C_{ij}=\int_{0}^{1}\frac{dy}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}}S_{i}K_{j},\\ S_{0,1,2}=\{-\ln X^{\prime},1-Y^{\prime}\ln X^{\prime},-\frac{1}{2}+Y^{\prime}-{Y^{\prime}}^{2}\ln X^{\prime}\},\\ K_{0,1,2,3,4,5}=\{1,y,1-y,y^{2},(1-y)^{2},y(1-y)\},\\ C_{5}=-\int_{0}^{1}dy\ln(y^{2}\mathrm{k}^{2}/\penalty 50m^{2}-y\mathrm{k}^{2}/m^{2}+1),\\ X^{\prime}=1+\frac{1}{Y^{\prime}},\;Y^{\prime}=\frac{m^{2}-y\mathrm{p}^{\prime 2}-(1-y)\mathrm{p}^{2}}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}},\end{gathered}$ (62)
and $\mathrm{k}=\mathrm{p}-\mathrm{p}^{\prime}$. In the Coulomb gauge, these coefficients can be determined from Ref. OP_CG:
@@
$\begin{gathered}A=\varepsilon_{a}^{2}(2F_{1}-F_{2})-F_{22}+m^{2}(2F_{1}-3F_{2})+p^{\prime 2}(F_{11}-F_{8}+4F_{5}-5F_{3})-\\ -4(\mathbf{p^{\prime}}\cdot\mathbf{p})^{2}F_{13}+p^{2}(F_{10}-F_{11}-5F_{2}+5F_{3}+4F_{4}-4F_{5}-F_{7}+\\ +F_{8}+(\mathbf{p^{\prime}}\cdot\mathbf{p})(-2F_{12}+4F_{13}))+(-4F_{5}+4F_{6}+2F_{8}-2F_{9})k^{2}+\\ +(\mathbf{p^{\prime}}\cdot\mathbf{p})(4F_{1}-2F_{19}-2F_{2}-2F_{14}k^{2})-\varepsilon_{a}^{2}(B_{1}+C_{1}-D).\end{gathered}$ (64)
The coefficients $F_{1-22}$ introduced above can be determined as follows. Let us define $B_{ij}^{k}$ as
$B_{i\,j}^{k}=\begin{cases}1&\text{if }i\leq k\leq j\\ 0&\text{otherwise}\end{cases}\,\,$, (65)
@@
$\begin{gathered}F_{i}=\int_{0}^{1}\frac{du}{\mathrm{t}^{2}}\Big[C_{1}^{i}\delta_{1}+C_{2}^{i}-C_{3}^{i}\frac{A}{\mathrm{t}^{2}}\Big]C_{4}^{i},\\ \delta_{1}=\ln\left(\frac{\mathrm{t}^{2}+A}{A}\right),\,\,C_{1}^{i}=[1-B_{2\,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)][1-B_{4\,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)],\\ C_{2}^{i}=B_{2\,3}^{i}+\frac{1}{2}B_{4\,6}^{i},\,\,C_{3}^{i}=B_{4\,6}^{i},\,\,C_{4}^{i}=\{1,1,u,1,u,u^{2}\}.\end{gathered}$ (66)
Here $A=m^{2}-u\mathrm{p}^{2}-(1-u)\mathrm{p}^{2}$, $\mathrm{t}=u\mathrm{p}^{\prime}+(1-u)\mathrm{p}$. For $F_{7-11}$ (below the coefficient $F_{7}$ is obtained by substituting $i=1$, and so on, i.e., $F_{8}\leftrightarrow i=2$, …):
$\begin{gathered}F_{i}=\int_{0}^{1}du\Big[C_{5}^{i}\delta_{2}-C_{6}^{i}\frac{1}{t^{2}}\Big]C_{7}^{i},\\ \delta_{2}=\frac{2}{\sqrt{t^{2}C}}\tanh^{-1}\left[\left(\frac{t^{2}}{C}\right)^{1/2}\right],\,\,C_{5}^{i}=\frac{C}{t^{2}}B_{1\,3}^{i}+B_{45}^{i},\\ C_{6}^{i}=2B_{1\,3}^{i},\,\,C_{7}^{i}=\{1,u,u^{2},1,u\},\end{gathered}$ (67)
where $C=m^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+t_{0}^{2}$. For $F_{12}-F_{21}$ (below $F_{12}\leftrightarrow i=1$):
$\begin{gathered}F_{i}=\int_{0}^{1}dsdu\Big[C_{8}^{i}\delta_{3}+C_{9}^{i}\delta_{4}\Big]C_{10}^{i},\\ \delta_{3}=\frac{1}{s(t^{2})^{2}}\left\{\left(\frac{B}{st^{2}}\right)^{1/2}\tanh^{-1}\Big[\left(\frac{st^{2}}{B}\right)^{1/2}\Big]-1\right\},\,\,\delta_{4}=\frac{1}{t^{2}(B-st^{2})},\\ C_{8}^{i}=-3B_{1\,7}^{i}+2t^{2}B_{8\,10}^{i},\,\,C_{9}^{i}=B_{1\,7}^{i},\\ C_{10}^{i}=\{1,u,u^{2},s,su,su^{2},su^{3},1,s,su\},\end{gathered}$ (68)
where $B=m^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+st_{0}^{2}$. And finally $F_{22}$ is given by
$F_{22}=\int_{0}^{1}du\ln(u^{2}\mathrm{k}^{2}/\penalty 50m^{2}-u\mathrm{k}^{2}/\penalty 50m^{2}+1).$ (69)
@@ Appendix C Many-potential term
Reduced matrix elements can be calculated using the PW expansion of the photon propagator. Let us illustrate this with the example of the Feynman gauge, where, considering the operator $I_{F}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ from Eq. (5), one can use the standard expression
$I_{F}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})\sim\frac{e^{i\omega r_{12}}}{r_{12}}=4\pi i\omega\sum_{L=0}^{\infty}\sum_{M=-L}^{L}j_{L}(\omega r_{<})h_{L}^{(1)}(\omega r_{>})Y_{LM}^{*}(\mathbf{\hat{r}_{1}})Y_{LM}(\mathbf{\hat{r}_{2}}),$ (70)
where $j_{L}$ and $h_{L}^{(1)}$ are the spherical Bessel and Hankel functions of the first kind respectively and $r_{>}=\max(r_{1},r_{2})$, $r_{<}=\min(r_{1},r_{2})$. Then, one can obtain
$\begin{gathered}\langle ab||I_{F}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}dr_{2}\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(\omega,r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}[J]g_{L}(\omega,r_{1},r_{2})D_{JL,ac}(r_{1})D_{JL,bd}(r_{2})\}\end{gathered}$ (71)
@@
$\begin{gathered}\langle ab||I_{C}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}dr_{2}\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(0,r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}a_{JL}g_{L}(\omega,r_{1},r_{2})D_{JL,ac}(r_{1})D_{JL,bd}(r_{2})\\ +(-1)^{J+1}b_{J}[g_{J}^{\text{ret}}(\omega,r_{1},r_{2})D_{JJ+1,ac}(r_{1})D_{JJ-1,bd}(r_{2})+g_{J}^{\text{ret}}(\omega,r_{2},r_{1})D_{JJ-1,ac}(r_{1})D_{JJ+1,bd}(r_{2})]\}\end{gathered}$ (72)
for the Coulomb gauge. In the above expressions, coefficients that include the radial parts of wave functions are:
$\begin{gathered}A_{ab}(r_{1})=G_{a}(r_{1})G_{b}(r_{1})+F_{a}(r_{1})F_{b}(r_{1}),\\ D_{JL,ab}(r_{1})=G_{a}(r_{1})F_{b}(r_{1})H_{L}^{J}(\kappa_{a},-\kappa_{b})-F_{a}(r_{1})G_{b}(r_{1})H_{L}^{J}(-\kappa_{a},\kappa_{b}),\\ G_{a}(r)=rg_{a}(r),\quad F_{a}(r)=rf_{a}(r).\end{gathered}$ (73)
The numerical coefficients are:
$\begin{gathered}a_{JL}=\begin{cases}J+1,&\text{for}\ L=J-1\\ 2J+1,&\text{for}\ L=J\\ J,&\text{for}\ L=J+1\end{cases}\;\;,\\ b_{J}=\sqrt{J(J+1)}\frac{\sqrt{[J+1][J-1]}}{[J]},\\ [a]=2a+1.\end{gathered}$ (74)
When considering the photon propagator, the function $g_{J}$ arises, which includes
$\begin{gathered}g_{J}(\omega,r_{1},r_{2})=i[J]\omega j_{J}(\omega r_{<})h_{J}^{(1)}(\omega r_{>}),\\ g_{J}(0,r_{1},r_{2})=\frac{r_{<}^{J}}{r_{<}^{J+1}}.\end{gathered}$ (75)
In the Coulomb gauge, we define the function $g_{J}^{\text{ret}}$:
$g_{J}^{\text{ret}}(\omega,r_{1},r_{2})=\begin{cases}i[J]\omega j_{J+1}(\omega r_{1})h_{J-1}^{(1)}(\omega r_{2}),&\text{for}\ r_{1}<r_{2}\\ i[J]\omega j_{J-1}(\omega r_{2})h_{J+1}^{(1)}(\omega r_{1})-\frac{[J]^{2}r_{2}^{J-1}}{\omega^{2}r_{1}^{J+2}},&\text{otherwise}\end{cases}\;\;.$ (76)
Angular coefficients are:
@@
Note that (72) is consistent with Ref. mel_yerokh, except for the typo in Eq. (B6) there.

### v2

@@ Appendix A Free-electron self-energy operator
$D_{\mu\nu}^{\text{F}}(\mathrm{k})=-\frac{g_{\mu\nu}}{\mathrm{k}^{2}}$ (52)
@@
$D_{\mu\nu}^{\text{C}}(\mathrm{k})=\frac{1}{\mathrm{k}^{2}}\left(-g_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{\mathbf{k}^{2}}+\frac{k^{0}(k_{\mu}\delta_{\nu 0}+k_{\nu}\delta_{\mu 0})}{\mathbf{k}^{2}}\right)$ (53)
in the Coulomb gauge. In Eq. (53), $\delta_{\nu\mu}$ is the four-dimensional Kronecker delta, $\delta_{\mu\nu}\equiv\delta^{\mu}_{\nu}$. The expression (51) is UV divergent in both gauges. To eliminate the UV divergences, we apply the renormalization procedure. To make sense of divergent expressions, we use the dimensional regularization $d=4-2\varepsilon$. In both gauges, we separate out the UV divergent part by writing the self-energy in the form:
@@
where $\Delta_{\varepsilon}=1/\varepsilon-\gamma_{E}+\ln 4\pi+\ln m_{e}$, $\gamma_{E}$ is the Euler constant, and the mass counterterm is
$\delta m=\frac{3\alpha}{4\pi}m_{e}\left(\Delta_{\varepsilon}+\frac{4}{3}\right).$ (55)
Note that following Ref. Yerokhin and Shabaev (1999) we use this definition of $\Sigma^{(0)}_{R}$ instead of writing a more familiar expression in terms of the renormalization constant $Z_{2}$. In the Feynman gauge, the second approach leads to the IR singularities of individual contributions. In the Coulomb gauge, both approaches are equally convenient, but, for consistency, we write the expression for the free-electron self-energy operator in this form.
@@
[…]
The coefficients $a$, $b$ and $c$ are gauge-dependent. In the Feynman gauge, they can be found, e.g., in Ref. Yerokhin and Shabaev (1999):
$\begin{gathered}a(p_{0},p)=2m_{e}\left(1+\frac{2\rho}{1-\rho}\ln\rho\right),\\ b(p_{0},p)=-\frac{2-\rho}{1-\rho}\left(1+\frac{\rho}{1-\rho}\ln\rho\right),\\ c(p_{0},p)=0,\end{gathered}$ (57)
where $\rho=1-{\mathrm{p^{2}}}/{m_{e}^{2}}$. For the Coulomb gauge, they could be derived from the expression obtained in Ref. Adkins (1983):
@@
$\begin{gathered}X=1+\frac{p^{2}}{m_{e}^{2}}(1-x),\\ Y=1-\frac{\mathrm{p}^{2}}{m_{e}^{2}}(1-x)-i0,\\ Z=1-\frac{p_{0}^{2}}{m_{e}^{2}}(1-u)+\frac{p^{2}}{m_{e}^{2}}(1-xu)-i0.\end{gathered}$ (59)
Some integrals in these expressions can be evaluated analytically, which was done, e.g., in Ref. Malenfant (1987). The latter results allows one to write for the coefficients $a$, $b$, and $c$:
$\begin{gathered}a(p_{0},p)=2m_{e}\left(1-F_{0}+\frac{\rho\log\rho}{1-\rho}\right),\\ b(p_{0},p)=\frac{(\rho-2)(1-\rho+\rho\ln\rho)}{(1-\rho)^{2}}-2F_{2}\rho+\frac{2m_{e}^{2}(F_{1}\rho\ln\rho-F_{0})}{p^{2}},\\ c(p_{0},p)=2\frac{p_{0}}{p^{2}}\big(F_{0}m_{e}^{2}-F_{1}m_{e}^{2}\rho\ln\rho+F_{2}\rho p^{2}\big),\end{gathered}$ (60)
where $F_{0}$, $F_{1}$, and $F_{2}$ are defined according to
$\begin{gathered}F_{0}=\frac{\sqrt{p^{2}+m_{e}^{2}}}{p}\ln\left|\frac{\sqrt{p^{2}+m_{e}^{2}}+p}{\sqrt{p^{2}+m_{e}^{2}}-p}\right|-2,\\ F_{1}=\frac{p_{0}}{p}\ln\left|\frac{p_{0}+p}{p_{0}-p}\right|-2,\\ F_{2}=\int_{0}^{1}dx\frac{\sqrt{x}\ln X}{X-\rho+i0}.\end{gathered}$ (61)
@@ Appendix B Free-electron vertex operator
In order to calculate the one-potential term, one needs the free-electron vertex function given by:
@@
This expression is UV divergent in both gauges. Using dimensional regularization, it can be written as:
@@
where, as in the case of the free-electron self-energy operator, the UV finite part of $\Gamma_{R}$ also turns out to be IR finite. For brevity, we discuss below only the time component of the operator, $\Gamma^{0}_{R}$, for the case when $p_{0}=p^{\prime}_{0}=\varepsilon_{a}$:
@@
$\begin{gathered}A=C_{5}-2+\mathrm{p}^{\prime 2}C_{11}+\mathrm{p}^{2}C_{12}+4(\mathrm{p}^{\prime}\cdot\mathrm{p})(C_{00}+C_{11}+C_{12})+m_{e}^{2}(-2C_{00}+C_{11}+C_{12}),\\ B_{1}=-4(C_{11}+C_{23}),\;B_{2}=-4(C_{00}+C_{11}+C_{12}+C_{25}),\\ C_{1}=B_{2},\;C_{2}=-4(C_{12}+C_{24}),\\ D=2(C_{00}+C_{11}+C_{12}),\\ H_{1}=4m_{e}(C_{00}+2C_{11}),\;H_{2}=4m_{e}(C_{00}+2C_{12}),\;G_{1}=G_{2}=0,\end{gathered}$ (65)
@@
$\begin{gathered}C_{ij}=\int_{0}^{1}\frac{dy}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}}S_{i}K_{j},\\ S_{0,1,2}=\{-\ln X^{\prime},1-Y^{\prime}\ln X^{\prime},-\frac{1}{2}+Y^{\prime}-{Y^{\prime}}^{2}\ln X^{\prime}\},\\ K_{0,1,2,3,4,5}=\{1,y,1-y,y^{2},(1-y)^{2},y(1-y)\},\\ C_{5}=-\int_{0}^{1}dy\ln(y^{2}\mathrm{k}^{2}/\penalty 50m_{e}^{2}-y\mathrm{k}^{2}/m_{e}^{2}+1),\\ X^{\prime}=1+\frac{1}{Y^{\prime}},\;Y^{\prime}=\frac{m_{e}^{2}-y\mathrm{p}^{\prime 2}-(1-y)\mathrm{p}^{2}}{(y\mathrm{p}^{\prime}+(1-y)\mathrm{p})^{2}},\end{gathered}$ (66)
and $\mathrm{k}=\mathrm{p}-\mathrm{p}^{\prime}$. For the Coulomb gauge, these coefficients can be readily determined from Ref. Holmberg (2011):
@@
$\begin{gathered}A=\varepsilon_{a}^{2}(2F_{1}-F_{2})-F_{22}+m_{e}^{2}(2F_{1}-3F_{2})+p^{\prime 2}(F_{11}-F_{8}+4F_{5}-5F_{3})-\\ -4(\mathbf{p^{\prime}}\cdot\mathbf{p})^{2}F_{13}+p^{2}(F_{10}-F_{11}-5F_{2}+5F_{3}+4F_{4}-4F_{5}-F_{7}+\\ +F_{8}+(\mathbf{p^{\prime}}\cdot\mathbf{p})(-2F_{12}+4F_{13}))+(-4F_{5}+4F_{6}+2F_{8}-2F_{9})k^{2}+\\ +(\mathbf{p^{\prime}}\cdot\mathbf{p})(4F_{1}-2F_{19}-2F_{2}-2F_{14}k^{2})-\varepsilon_{a}^{2}(B_{1}+C_{1}-D).\end{gathered}$ (68)
The coefficients $F_{1-22}$ introduced above can be determined as follows. Let us define $B_{i,j}^{k}$ as
$B_{i,j}^{k}=\begin{cases}1\,,&\text{if }i\leq k\leq j\,,\\ 0\,,&\text{otherwise}\,.\end{cases}$ (69)
@@
$\begin{gathered}F_{i}=\int_{0}^{1}\frac{du}{\mathrm{t}^{2}}\Big[C_{1}^{i}\delta_{1}+C_{2}^{i}-C_{3}^{i}\frac{A}{\mathrm{t}^{2}}\Big]C_{4}^{i},\\ \delta_{1}=\ln\left(\frac{\mathrm{t}^{2}+A}{A}\right),\,\,C_{1}^{i}=[1-B_{2,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)][1-B_{4,6}^{i}\left(1+\frac{A}{\mathrm{t}^{2}}\right)],\\ C_{2}^{i}=B_{2,3}^{i}+\frac{1}{2}B_{4,6}^{i},\,\,C_{3}^{i}=B_{4,6}^{i},\,\,C_{4}^{i}=\{1,1,u,1,u,u^{2}\}.\end{gathered}$ (70)
Here, $A=m_{e}^{2}-u\mathrm{p}^{2}-(1-u)\mathrm{p}^{2}$, $\mathrm{t}=u\mathrm{p}^{\prime}+(1-u)\mathrm{p}$. For $F_{7-11}$ (below the coefficient $F_{7}$ is obtained by substituting $i=1$, and so on, i.e., $F_{8}\leftrightarrow i=2$, …):
$\begin{gathered}F_{i}=\int_{0}^{1}du\Big[C_{5}^{i}\delta_{2}-C_{6}^{i}\frac{1}{t^{2}}\Big]C_{7}^{i},\\ \delta_{2}=\frac{2}{\sqrt{t^{2}C}}\tanh^{-1}\left[\left(\frac{t^{2}}{C}\right)^{1/2}\right],\,\,C_{5}^{i}=\frac{C}{t^{2}}B_{1,3}^{i}+B_{4,5}^{i},\\ C_{6}^{i}=2B_{1,3}^{i},\,\,C_{7}^{i}=\{1,u,u^{2},1,u\},\end{gathered}$ (71)
where $C=m_{e}^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+t_{0}^{2}$. For $F_{12}-F_{21}$ (below $F_{12}\leftrightarrow i=1$):
$\begin{gathered}F_{i}=\int_{0}^{1}dsdu\Big[C_{8}^{i}\delta_{3}+C_{9}^{i}\delta_{4}\Big]C_{10}^{i},\\ \delta_{3}=\frac{1}{s(t^{2})^{2}}\left\{\left(\frac{B}{st^{2}}\right)^{1/2}\tanh^{-1}\Big[\left(\frac{st^{2}}{B}\right)^{1/2}\Big]-1\right\},\,\,\delta_{4}=\frac{1}{t^{2}(B-st^{2})},\\ C_{8}^{i}=-3B_{1,7}^{i}+2t^{2}B_{8,10}^{i},\,\,C_{9}^{i}=B_{1,7}^{i},\\ C_{10}^{i}=\{1,u,u^{2},s,su,su^{2},su^{3},1,s,su\},\end{gathered}$ (72)
where $B=m_{e}^{2}-u\mathrm{p^{\prime}}^{2}-(1-u)\mathrm{p}^{2}+st_{0}^{2}$. And, finally, $F_{22}$ is given by
$F_{22}=\int_{0}^{1}du\ln(u^{2}\mathrm{k}^{2}/\penalty 50m_{e}^{2}-u\mathrm{k}^{2}/\penalty 50m_{e}^{2}+1).$ (73)
@@ Appendix C Many-potential term
Reduced matrix elements of the operators $I_{\text{F}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ and $I_{\text{C}}(\omega,\mathbf{r}_{1},\mathbf{r}_{2})$ can be derived using the PW expansion of the photon propagator. Namely, for this aim one needs the standard expansion formulas Varshalovich et al. (1988) for the functions which depend on $r_{12}$, such as
$\frac{e^{i\omega r_{12}}}{r_{12}}=4\pi i\omega\sum_{L=0}^{\infty}\sum_{M=-L}^{L}j_{L}(\omega r_{<})h_{L}^{(1)}(\omega r_{>})Y_{LM}^{*}(\mathbf{\hat{r}_{1}})Y_{LM}(\mathbf{\hat{r}_{2}}).$ (74)
Here, $j_{L}$ and $h_{L}^{(1)}$ are the spherical Bessel function and spherical Hankel functions of the first kind, respectively and $r_{>}=\max(r_{1},r_{2})$ and $r_{<}=\min(r_{1},r_{2})$. Applying these formulas and performing the integrations over the angular variables, after some tedious manipulations one can obtain
$\begin{gathered}\langle ab||I_{\text{F}}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}r_{1}^{2}\int_{0}^{\infty}dr_{2}r_{2}^{2}\,\Big\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(\hat{\omega},r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}[J]g_{L}(\hat{\omega},r_{1},r_{2})D^{JL}_{ac}(r_{1})D^{JL}_{bd}(r_{2})\Big\}\end{gathered}$ (75)
@@
$\begin{gathered}\langle ab||I_{\text{C}}(\omega)||cd\rangle_{J}=\alpha\int_{0}^{\infty}dr_{1}r_{1}^{2}\int_{0}^{\infty}dr_{2}r_{2}^{2}\,\Big\{(-1)^{J}G_{J}(\kappa_{a},\kappa_{c})G_{J}(\kappa_{b},\kappa_{d})g_{J}(0,r_{1},r_{2})A_{ac}(r_{1})A_{bd}(r_{2})\\ +\sum_{L}(-1)^{L+1}a_{JL}g_{L}(\hat{\omega},r_{1},r_{2})D^{JL}_{ac}(r_{1})D^{JL}_{bd}(r_{2})\\ +(-1)^{J+1}b_{J}[g_{J}^{\text{ret}}(\hat{\omega},r_{1},r_{2})D^{JJ+1}_{ac}(r_{1})D^{JJ-1}_{bd}(r_{2})+g_{J}^{\text{ret}}(\hat{\omega},r_{2},r_{1})D^{JJ-1}_{ac}(r_{1})D^{JJ+1}_{bd}(r_{2})]\Big\}\end{gathered}$ (76)
for the Coulomb gauge. In the above expressions, $[J]=2J+1$ and the following functions are defined:
$g_{J}(\omega,r_{1},r_{2})$ $=i[J]\omega j_{J}({\omega}r_{<})h_{J}^{(1)}({\omega}r_{>}),$ (77)
$g_{J}(0,r_{1},r_{2})$ $=\frac{r_{<}^{J}}{r_{<}^{J+1}},$ (78)
@@
The expressions depending on the radial wave functions read as
$A_{ab}(r)$ $=g_{a}(r)g_{b}(r)+f_{a}(r)f_{b}(r)\,$, (80)
$D^{JL}_{ab}(r)$ $=g_{a}(r)f_{b}(r)H_{L}^{J}(\kappa_{a},-\kappa_{b})-f_{a}(r)g_{b}(r)H_{L}^{J}(-\kappa_{a},\kappa_{b})\,.$ (81)
Finally, the angular coefficients are
[…]
$H_{L}^{J}(\kappa_{a},\kappa_{b})$ $=(-1)^{l_{a}}\sqrt{6[j_{a}][j_{b}][l_{a}][l_{b}]}{\begin{pmatrix}l_{a}&L&l_{b}\\ 0&0&0\end{pmatrix}}{\begin{Bmatrix}j_{a}&\frac{1}{2}&l_{a}\\ J&1&L\\ j_{b}&\frac{1}{2}&l_{b}\end{Bmatrix}},$ (83)
$a_{JL}$ $=\begin{cases}J+1,&\text{for}\ L=J-1,\\ 2J+1,&\text{for}\ L=J,\\ J,&\text{for}\ L=J+1,\end{cases}$ (84)
$b_{J}$ $=\sqrt{J(J+1)}\frac{\sqrt{[J+1][J-1]}}{[J]}.$ (85)
Note that (76) is consistent with Ref. Yerokhin et al. (2021), except for the typo in Eq. (B6) there.